OpenStudy (anonymous):

how do i differentiate this using implicit differentiation y^2=9x^2+8x

OpenStudy (anonymous):

2y?

OpenStudy (loser66):

@satellite73 I am not good at explanation, please help her

OpenStudy (anonymous):

so do i need to solve it for y first?

OpenStudy (anonymous):

@zepdrix @dumbcow @Directrix

zepdrix (zepdrix):

Oh sorry I musta missed this one :x you figure it out already?

OpenStudy (dumbcow):

differentiate everything with respect to "x" "y" is a function of "x" ...y(x) so by chain rule when you differentiate the "y" term, you take derivative wrt "y" and multiply by dy/dx \[\rightarrow 2y \frac{dy}{dx} = 18x+8\] solve for dy/dx \[\frac{dy}{dx} = \frac{9x+4}{y}\]